# Conditional Balance: Improving Multi-Conditioning Trade-Offs in Image Generation

### v1

@@ 1 Introduction
In this work, we investigate the style-content trade-offs of different conditional inputs seperately and in combination in generative models. Our experiments reveal that many issues in conditional generation arise from over-conditioning and the combination of conditionals that were underrepresented during model training. We analyze a diffusion generation process (using SDXL’s architecture [21]) and isolate different aspects of its generative capabilities.
@@
Our complex-conditioning evaluation set and code are available through our project page: conditional-balance.github.io/
@@ 2 Related Work
DDPMs [6, 25, 23, 28, 21] achieved a significant leap in generating novel images conditioned on a text prompts by refining over text-image paired datasets [29]. While these models combine content and style effectively, their sole reliance on text conditioning comes with shortcomings, such as failure to align to complex text conditions [20, 34, 37] and an inferior ability to generate content and style combinations which were underrepresented during training.
To overcome these limitations various Conditioning methods were developed, influencing the generated image to resemble information from another image. ControlNet [38] conditions the output with structure information using image maps such as Canny, depth, and pose. Gal et al. [9], and others [10, 27, 1], condition the generated image to preserve unique (personalized) properties of an object.
Image-based style conditioning is a thoroughly researched area [36, 8, 33, 16, 13, 4, 31, 5, 26], usually demonstrated on SDXL [21] for its artistic superiority. IP-Adapter [36] and InstantStyle [33] inject the style using a pretrained dedicated Cross-Attention layer [32] for the conditioning image. B-LoRA [8] and ZipLoRA [30] uses LoRA [14] to fine-tune residual weights to match content and style of conditioning images. StyleAligned method [13] generates a style image in parallel while generating the output image and injects the style information using AdaIN [15] between Self-Attention layers [32] while Jeong et al. [16] extend this idea by replacing the attention feature directly.
@@
Recent works [39, 8, 16, 33] analyze the diffusion process to pinpoint parts responsible for generating various visual aspects. ProSpect [39] examine how varying text conditions at different timesteps impact aspects like material, artistic style, and content alignment. Our findings align with theirs, revealing that each timestep serves a unique conditioning function. However, we extend this analysis by investigating the model’s internal layers, allowing for the use of multiple conditional inputs without interference. While ProSpect focuses exclusively on text conditioning, our study incorporates both text and image conditioning. Style conditioning works [8, 33], examine internal layers to pinpoint those sensitive to content and style. B-LoRA tests prompt stylization effects in different attention blocks and employs LoRA [14] to fine-tune style sensitive weights for specific images. InstantStyle analyzes SDXL attention layers for content alignment and style sensitivity and concludes the same style-sensitive blocks. In contrast, our approach thoroughly evaluates each self-attention layer independently, ranking its sensitivity for each timestep. This methodology maximizes style and content conditioning while minimizing interference between layers, thus enhancing stylistic freedom and reducing content artifacts.
@@ 3 Content-Style Tradeoff
In the following section, we expand our approach, which includes a novel method for analyzing style sensitive layers in SDXL, and a way to apply this knowledge to balance the use of conditionals at inference time without the need for any additional training.
@@ 5.1 Style Sensitivity
Figure 5: Collection Example. Examples from the collections we use for sensitivity analysis. The object depicted in the collection is constant (cat), where each cluster holds a different style for style conditioning (top). At the bottom, we use “black and white ink” style to normalize color balance for structure conditioning.
@@ 5.2 Structure Sensitivity
Applying our analysis method to this collection shows sensitivity in various Up layers at all timesteps. Reducing ControlNet input in these layers reveals geometric style changes as more timesteps are used. We conducted an ablation study (see supplemental file) to better understand these effects and discovered that the Up layers seem to handle the fine details from control maps, while Middle layers maintain overall structure alignment. Therefore, using our analysis we can rank the layers and timesteps according to structure sensitivity and apply conditioing (this time controlNet input) only to a subset of the layers, allowing more balanced control with other conditions. Note that this kind of geometric style control cannot be achieved by simply reducing the scale parameter of ControlNet (see Fig. 4).
@@ 6 Results
Instead of choosing the number of layers K, we define two interpolation parameters: $\lambda_{S}$ and $\lambda_{T}$, that control the percentage of layers used for style and structure respectively. This allows users to control different content-style balancing ratio, so we encourage tuning $\lambda_{S}$ and $\lambda_{T}$. However, to be robust in our comparisons we set the layer subset size to 30 ($\lambda_{S}=0.43$) for both text and content image conditioned generations and limit the content control over 850 timesteps ($\lambda_{T}=0.15$) which induces geometric style and freedom. We find these values robust for various styles. We present results in Fig. 8 and in the supplemental file.
@@ 6.1 Evaluation Details
Figure 9: Style Transfer. A sample of style transfer using B-LoRA with an ungenerated style image (top row) and a content image (left column). Our balancing stratagy improves style alignment and consistency for different content inputs.
6.2 Comparisons
We apply our balancing strategy to StyleAligned and B-LoRA and compare our method to four style-based approaches: StyleAligned and B-LoRA (without balancing), InstantStyle, and Jeong et al. (we leave their pipeline analysis and balancing for future work) . We also compare to other methods in the supplemental file. As B-LoRA pre-trains residual LoRA weights prior to inference, we replace the trained layers to the ones analyzed by our method. For fairness, we base our decision on the same number of attention layers (20) used by B-LoRA, although we find that in some scenarios using a larger number of layers improves results.
@@ 6.2 Comparisons
We report qualitative and quantitative results in Fig. 6 and Fig. 7, respectively. As demonstrated in Fig. 6 our layer balancing strategy preserving both style and structure (right columns) and is more cosistent over various conditioning combinations. Meanwhile, other methods shows inconsistent output quality, where some combinations show satisfying results while others show apparent issues such as general and geometric style loss, and content artifacts.
This balancing effect is also apparent quantitatively in Fig. 7. While other methods show varying content and style quality on different conditioning combinations (triangle vs. rhombus), balanced methods show consistent high scores in both style and content similarity.
@@
1. Multi-choice 386 (210) 244 (420) 630 35.1 $0.001$
2. A/B B-LoRA 195 (126) 57 (126) 252 30.0 $0.001$
3. A/B StyleAligned 185 (126) 67 (126) 252 21.9 $0.001$
@@ 6.3 User Study
6.4 Style Transfer
We experiment with additional methods such as style transfer and material style generation. We use the balanced version of B-LoRA, which base their style conditioning on non-invertive methods, which in our experience were found less optimal. We present a sample of our qualitative results in Fig. 14. Additional results and material style transfer results can be found in the supplemental file.
@@ 7 Discussion and Limitations
In this paper we investigated the tradeoff between content and style conditionals and showed that over-conditionin prevents content and style aligment. We further present a novel method for analyzing sensitivity in self-attention layers of DDPMs by measuring the clustering score of the representation of a collection of images. We use this method to balance multi conditions by limiting conditioning on sensitive layers only, demonstrating improved results.
Our analysis was limited to the SDXL architecture, as it underlies most recent stylization methods. Additionally, since our current approach focuses on self-attention layers, stylization techniques relying on different architectural components may require separate evaluation. We leave these extensions to future work.
@@ 8.1 Evaluation Set
In Sec. 6 we extend our evaluation set with 22 additional styles and 10 additional prompts. The additional styles are: Pixar, Pixel Art, Edvard Munch, Franz Marc, John James Audubon, Oswaldo Guayasamin, Henri Matisse, Wassily Kandinsky, Ilya Repin, Gustav Klimt, Voxel Art, Vector Art, Anime, Henri De Toulouse-Lautrec, Yoshitaka Amano, Cyberpunk, Concept Art, Low Poly, Gustav Courbet, Paul Cézanne, Jean Metzinger and Georges Seurat. The additional prompt are: “An old TV set,” “A colorful fishbowl,” “A house in a village,” “A bartender leaning on his bar,” “A brown horse galloping” (Easy) and “A robot wearing a fedora holding a flower,” “A humpback whale floating in the sky carried by large colorful balloons,” “A fantasty castle with blue pointy rooftops located on a hill in a green valley,” “An orc and a blond wood-elf sitting in a tavern drinking beer as friends,” “Gandalf the Gray riding a horse while casting a spell with his wooden staff” (Complex).
@@ 9.2 Layer Rankings
We present an example of a a choice of 30 layers of Key layers for both style and geometry in Fig. 12. As can be seen the majority of layers show consistency over time while some layers change on various timesteps.
9.3 Geometric Style Ablation Study
@@
10 Appendix C. — Results
10.1 Qualtitative Results.
To demonstrate the ability of our conditioning strategy we show additional results produced by the Balanced versions of StyleAligned and B-LoRA for “Text Only” (Fig. 15), “Text + Canny” (Fig. 16), and “Text + Depth”’ (Fig. 17). We share additional qualitative comparisons between the balanced versions of StyleAligned and B-LoRA with the benchmark methods from Sec. 6 in Figs. 18 to 21. In addition, we compare ourself to two additional recent methods: RB-Modulation [26] and InstaStyle [5]. Since both methods provide access to their model only through an interactive web interface we could not evaluate their results using Canny or Depth conditioning. For this reason we refrained from using their methods in our main comparison in Sec. 6. For fairness, we present a qualitative comparison in Fig. 22 using a style reference and a text prompt without any structure conditioning.

### v2

@@ 1 Introduction
In this work, we investigate the style-content trade-offs of different conditional inputs separately and in combination in generative models. Our experiments reveal that many issues in conditional generation arise from over-conditioning and the combination of conditionals that were underrepresented during model training. We analyze a diffusion generation process (using SDXL’s architecture [26]) and isolate different aspects of its generative capabilities.
@@
Our complex-conditioning evaluation set and code are available through our project page.
@@ 2 Related Work
DDPMs [9, 30, 28, 33, 26] achieved a significant leap in generating novel images conditioned on text prompts by refining over text-image paired datasets [34]. While these models combine content and style effectively, their sole reliance on text conditioning comes with shortcomings, such as failure to align to complex text conditions [24, 39, 42] and an inferior ability to generate content and style combinations which were underrepresented during training.
To overcome these limitations, various conditioning methods were developed, influencing the generated image to resemble information from another image. ControlNet [43] conditions the output with structure information using image maps such as Canny, depth, and pose. Gal et al. [13], and others [14, 32, 2], condition the generated image to preserve unique (personalized) properties of an object.
Image-based style conditioning is a thoroughly researched area [41, 12, 38, 20, 17, 6, 36, 8, 31], usually demonstrated on SDXL [26] for its artistic superiority. IP-Adapter [41] and InstantStyle [38] inject the style using a pretrained dedicated Cross-Attention layer [37] for the conditioning image. B-LoRA [12] and ZipLoRA [35] uses LoRA [18] to fine-tune residual weights to match content and style of conditioning images. StyleAligned method [17] generates a style image in parallel to generating the output image and injects the style information using AdaIN [19] between Self-Attention layers [37] while Jeong et al. [20] extend this idea by replacing the attention feature directly.
@@
Recent works [44, 12, 20, 38] analyze the diffusion process to pinpoint parts responsible for generating various visual aspects. ProSpect [44] examine how varying text conditions at different timesteps impact aspects like material, artistic style, and content alignment. Our findings align with theirs, revealing that each timestep serves a unique conditioning function. However, we extend this analysis by investigating the model’s internal layers, allowing for the use of multiple conditional inputs without interference. While ProSpect focuses exclusively on text conditioning, our study incorporates both text and image conditioning. Style conditioning works [12, 38], examine internal layers to pinpoint those sensitive to content and style. B-LoRA tests prompt stylization effects in different attention blocks and employs LoRA [18] to fine-tune style sensitive weights for specific images. InstantStyle analyzes SDXL attention layers for content alignment and style sensitivity and concludes the same style-sensitive blocks. In contrast, our approach thoroughly evaluates each self-attention layer independently, ranking its sensitivity for each timestep. Additionally, our method can analyze other attention architectures, like the Joint-Attention layer used by the more recent SD3 model family [11]. This methodology maximizes style and content conditioning while minimizing interference between layers, thus enhancing stylistic freedom and reducing content artifacts.
@@ 3 Content-Style Tradeoff
In the following section, we expand our approach, which includes a novel method for analyzing style sensitive layers in diffusion models, and a way to apply this knowledge to balance the use of conditionals at inference time without the need for any additional training.
@@ 5 Analysis
We conduct a similar analysis on the SD3.5-Large architecture [11]. Details can be found in the supplemental file.
@@ 5.1 Style Sensitivity
Figure 5: Collection Example. Collection examples used for structure sensitivity analysis. The object depicted in the collection is constant (cat), where each cluster holds a different style for style conditioning (top). At the bottom, we use “black and white ink” style to normalize color balance for structure conditioning.
@@ 5.2 Structure Sensitivity
Applying our analysis method to this collection shows sensitivity in various Up layers at all timesteps. Reducing ControlNet input in these layers reveals geometric style changes as more timesteps are used. We conducted an ablation study (see supplemental file) to better understand these effects and discovered that the Up layers seem to handle the fine details from control maps, while Middle layers maintain overall structure alignment. Therefore, using our analysis we can rank the layers and timesteps according to structure sensitivity and apply conditioning (this time ControlNet input) only to a subset of the layers, allowing more balanced control with other conditions. Note that this kind of geometric style control cannot be achieved by simply reducing the scale parameter of ControlNet (see Fig. 4).
@@ 6 Results
Instead of choosing the number of layers K, we define two interpolation parameters: $\lambda_{S}$ and $\lambda_{T}$, that control the percentage of layers used for style and structure respectively. This allows users to control different content-style balancing ratio, so we encourage tuning $\lambda_{S}$ and $\lambda_{T}$. However, to be robust in our comparisons we set the layer subset size to 30 ($\lambda_{S}=0.43$) for both text and content image conditioned generations and limit the content control over 850 timesteps ($\lambda_{T}=0.15$) which induces geometric style and freedom. We find these values robust for various styles, including non-painting ones (see supplemental file.) We present results in Fig. 6 and in the supplemental file.
@@ 6.1 Evaluation Details
6.2 Comparisons
We apply our balancing strategy to StyleAligned and B-LoRA and compare our method to four style-based approaches: StyleAligned and B-LoRA (without balancing), InstantStyle, and Jeong et al. (we leave their pipeline analysis and balancing for future work) . We also compare to other methods [31, 8] in the supplemental file. As B-LoRA pre-trains residual LoRA weights prior to inference, we replace the trained layers to the ones analyzed by our method. For fairness, we base our decision on the same number of self-attention layers (10) used by B-LoRA, although we find that in some scenarios using a larger number of layers improves results.
@@ 6.2 Comparisons
We report qualitative and quantitative results in Fig. 7 and Fig. 8, respectively. As demonstrated in Fig. 7 our layer balancing strategy preserving both style and structure (right columns) and is more consistent over various conditioning combinations. Meanwhile, other methods shows inconsistent output quality, where some combinations show satisfying results while others show apparent issues such as general and geometric style loss, and content artifacts.
This balancing effect is also apparent quantitatively in Fig. 8. While other methods show varying content and style quality on different conditioning combinations (triangle vs. rhombus), balanced methods show consistent high scores in both style and content similarity. We further evaluate these results using an additional style-representation using GRAM matrices [15] (See supplemental file.)
@@
1. Multi-choice 386 (210) 244 (420) 630 35.1 0.0010.0010.001
2. A/B B-LoRA 195 (126) 57 (126) 252 30.0 0.0010.0010.001
3. A/B StyleAligned 185 (126) 67 (126) 252 21.9 0.0010.0010.001
@@ 6.3 User Study
Figure 9: Style Transfer. A sample of style transfer using B-LoRA with an ungenerated style image (top row) and a content image (left column). Our balancing strategy improves style alignment and consistency for different content inputs.
6.4 Additional Applications
We experiment with additional applications such as style transfer [15, 19, 6, 22], material-style generation [7], and flexibly-conditioned content editing [25, 1, 3]. We present style transfer results using balanced B-LoRA in Fig. 20. Additional results and implementation details for all applications can be found in the supplemental file.
@@ 7 Discussion and Limitations
In this paper we investigated the tradeoff between content and style conditionals and showed that over-conditioning prevents content and style alignment. We further present a novel method for analyzing sensitivity in attention layers of DDPMs by measuring the clustering score of the representation of a collection of images. We use this method to balance multi conditions by limiting conditioning on sensitive layers only, demonstrating improved results.
We found that the main limitations of our method arise from its dependence on the capabilities of the base model. For instance, when generating images with a style unfamiliar to the base model, the result may exhibit an unintended style, as the model lacks sufficient knowledge to properly generate the style reference, leading to a style mismatch. We demonstrate these limitations in the supplemental file.
Acknowledgment
This work was partly supported by Joint NSFC-ISF Research Grant no. 3077/23
@@ 8.1 Evaluation Set
In Sec. 6 we extend our evaluation set with 22 additional styles and 10 additional prompts. The additional styles are: Pixar, Pixel Art, Edvard Munch, Franz Marc, John James Audubon, Oswaldo Guayasamin, Henri Matisse, Wassily Kandinsky, Ilya Repin, Gustav Klimt, Voxel Art, Vector Art, Anime, Henri De Toulouse-Lautrec, Yoshitaka Amano, Cyberpunk, Concept Art, Low Poly, Gustav Courbet, Paul Cézanne, Jean Metzinger and Georges Seurat. The additional prompt are: “An old TV set,” “A colorful fishbowl,” “A house in a village,” “A bartender leaning on his bar,” “A brown horse galloping” (Easy) and “A robot wearing a fedora holding a flower,” “A humpback whale floating in the sky carried by large colorful balloons,” “A fantasy castle with blue pointy rooftops located on a hill in a green valley,” “An orc and a blond wood-elf sitting in a tavern drinking beer as friends,” “Gandalf the Gray riding a horse while casting a spell with his wooden staff” (Complex).
@@ 9.2 Layer Rankings
We present an example of a choice of 30 layers of Key layers for both style and geometry in Fig. 12. As can be seen the majority of layers show consistency over time while some layers change on various timesteps.
Figure 13: Style Layers Ablation. Ablation study for the ability of our sensitivity analysis method. We investigate and compare the method’s ability to identify both style-sensitive and style-insensitive layers by using both layer rankings for stylization, where using insensitive layers is marked in blue and using sensitive layers is marked in green. We show quantitative results (top) and compare both scenarios to a random layers selection (red) and show a visual example for demonstrative purposes (bottom).
9.3 Style Layers Ablation Study - $\lambda_{S}$
In Sec. 3 of the main manuscript, we examine the trade-off between style and content in conditional image generation and provide an initial evaluation of the effectiveness of applying stylization to a subset of self-attention layers, as determined by our analysis method. In this subsection, we further explore our method’s ability to identify style sensitivities by assessing the performance of layers marked as not style-sensitive. We compare their impact to both random selection and our previous results from Sec. 3. To achieve this, we generate the evaluation set described in Sec. 3 using different subset sizes of style layers. However, instead of applying stylization to the $k$-most style-sensitive layers, we now apply it to the $k$-most insensitive layers.
Our experimental findings are illustrated in Fig. 13. Similar to the approach in Sec. 3 of the main manuscript, we provide quantitative evaluations for both style similarity (top left) and content similarity (top right), alongside a qualitative example for visual demonstration.
As shown in the top left plot, applying stylization to layers identified as not style-sensitive (blue graph) results in a significant reduction in the style similarity of the generated images compared to stylization using the style-sensitive layers (green graph). Furthermore, applying stylization to these layers leads to lower style similarity than random layer selection, reinforcing the effectiveness of our method in correctly identifying both style-sensitive and style-insensitive layers. Notably, these insensitive layers appear to be entirely unrelated to style, as stylizing them negatively impacts the overall stylization quality. Additionally, using the insensitive layers causes a slower increase in style similarity, which only starts to improve around (E.), when the style-sensitive layers take effect. Observing the content similarity plot (top-right), we can confirm our hypothesis that injecting style information into style-insensitive layers (blue) is not only ineffective for style similarity but also degrades content similarity. This degradation results in lower content similarity compared to both style-sensitive layers (green) and randomly selected layers (red.)
The qualitative impact of these findings is visually demonstrated in the image sequence at the bottom of Fig. 13. This sequence compares the interpolation effect of applying $\lambda_{S}$ on style-sensitive layers (top row) versus style-insensitive layers (bottom row). Each column ($A.-F.$), corresponding to points in the quantitative graphs, represents an increasing $\lambda_{S}$ value. Columns (A) and (F) show generated images for (A) no image style conditioning and (F) full layer style conditioning, both presenting a suboptimal result. The following observations can be made: (B) When using only 10 style-sensitive layers, the generated image already exhibits strong style representation while maintaining content integrity, whereas using style-insensitive layers results in no noticeable stylization effect. (C) Utilizing 14 sensitive layers maintains the previous quality, whereas using insensitive layers introduces content artifacts without achieving style alignment, likely due to injecting style information into content-related layers. (D) Applying 30 sensitive layers further improves both style and content alignment, while using insensitive layers results in some style improvements but also introduces content artifacts. (E) With 45 layers, stylization using our identified sensitive layers begins to introduce content artifacts, while using insensitive layers finally achieves reasonable style alignment as style and content-related layers start to blend in both cases.
Figure 15: SD3.5-Large Results. Results for text-only stylization, full layer stylization and balanced conditioning with SD3.5-Large. Prompts: “A black bear riding a bicycle in bustling market. The market is full of stands selling fruits and vegetables,” “a Siamese cat wearing scuba diving gear, horizontally scuba diving in a deep blue sea, watching a school of colorful jellyfish,” “A woman wearing a dark orange trench coat and large sunglasses walking in the cold streets of London,” “A penguin riding a motorcycle,” “A pirate ship sailing in the ocean and being attacked by the Kraken,” “A light-blue haired woman wearing a black attire and black steampunk goggles, leaning on a futuristic yellow motorcycle,” “An elephant painting the Savannah. He is sitting in front of a canvas holding a paint brush with his trunk,” and “A sculptor working in his studio. He is in the middle of sculpting a marble statue which starts to resemble a female figure.”
9.4 Geometric Style Ablation Study - $\lambda_{T}$
@@
Figure 16: Collection Examples. Representatives from the style clusters generated during the SD3.5-Large analysis. Each sample represents a different style.
@@
10 Appendix C. — Stable Diffusion 3
10.1 Style Conditioning
Recently, the Stable Diffusion 3 (SD3) model family [11] was released, offering new image generation diffusion models. Unlike SDXL [26] which uses a UNet based on Self-Attention and Cross-Attention layers, the SD3 models are based on a Joint-Attention layers which processes both image information and text information. For this reason to apply balanced conditioning on these models we adapted the stylization ideas suggested by Hertz et al. [17], which are Self-Attention based, to the Join-Attention architecture.
Like Hertz et al. [17], we apply stylization by applying AdaIN [19] between the attention features of a generated style image and the target image, and sharing the features of the Keys and Values on their projections. Since the Joint-Attention layer concatenates the Query, Key, and Value projections to the text encodings, we apply these operations before the concatenation to prevent changes in the text features. During our experiments we noticed that unlike Hertz et al. which applies AdaIN only on the Key and Query projections of the attention layers, applying AdaIN on the Value projections significantly contributes to the stylization of the target image and is key for achieving a satisfying result. For this reason we incorporate this change to our style conditioning algorithm for SD3.5-Large.
Figure 17: Gram Based Evaluation. Style and content evaluation using Gram-Matrix representation and Clip embeddings.
10.2 Analysis
We use our method presented in Sec. 4 to find SD3.5-Large style sensitivities as presented in Sec. 5.1. Fig. 16 shows examples from the collections generated during the analysis process where each example is taken from a different style cluster.
As shown in and Fig. 15, SD3.5-Large, like SDXL, struggles with complex conditioning combinations. This is particularly evident in the ”Text-Only Stylization” columns, where including only the artist’s name in the prompt often results in style mismatches and, in extreme cases, the complete omission of the target style. Conditioning on all layers for style leads to significant content drift, whereas our balancing method effectively aligns the generated results with both the content and style of the reference image. The balanced results we present were achieved by applying style conditioning to 28 out of 38 joint-attention layers ($\lambda_{S}=0.73$), which we identified as the optimal setting.
Figure 18: Photographic-Editing Styles. ”A tabby cat”. Style sensitive layers analyzed by our method are not limited to painting styles, but show sensitivities to other styles like photographic editing styles. Notice that using full conditioning may result with content drift and artifacts.
11 Appendix D. — Results
11.1 Qualitative Results
To demonstrate the ability of our conditioning strategy we show additional results produced by the Balanced versions of StyleAligned and B-LoRA for “Text Only” (Fig. 21), “Text + Canny” (Fig. 22), and “Text + Depth”’ (Fig. 23). We share additional qualitative comparisons between the balanced versions of StyleAligned and B-LoRA with the benchmark methods from Sec. 6 in Figs. 24 to 27. In addition, we compare ourself to two additional recent methods: RB-Modulation [31] and InstaStyle [8]. Since both methods provide access to their model only through an interactive web interface we could not evaluate their results using Canny or Depth conditioning. For this reason we refrained from using their methods in our main comparison in Sec. 6. For fairness, we present a qualitative comparison in Fig. 28 using a style reference and a text prompt without any structure conditioning.
11.2 Style Variation
To assess the generality of our method in identifying style-sensitive layers, we extend our experiments to styles beyond artistic paintings. Fig. 18 illustrates the applicability of our style layers to photographic-editing styles. Notably, the same style layers identified for artistic paintings enable our method to generate photographs that adopt a reference style while preserving content integrity and avoiding unwanted artifacts and content drift from the style image.
Figure 19: Unfamiliar Styles. A limitation example, where the style of an artist (A) is unknown to the base model, causing a mismatch in the reference image (B) which then results with a style mismatch in the target image (C), even when the content matches the artist.